i really need help solving this problem ive aldready done a few steps but not sure if im doing them right.
A 10% acid solution is to be mixed with a 50% solution to get 200 liters of a 20% solution. How many liters of each solution should be used ?

Answer by mananth(16949)

percent ----------------quantity
Acid solution A 10 ----------------x liters
Acid solution B 50 ----------------200-x liters
Total 20.00% ---------------- 200
..............
10x+50(200-x)=20*200
10x+10000-50x=4000
10x-50x =4000-10000
-40x=-6000
/-40
x=150 liters 10.00%
50 liters 50.00%